OpenStudy (anonymous):

Obtain a result for calculating the time required to consume a finite resource. dM/dt=-Pexp(at)+Fexp(-ut)

OpenStudy (anonymous):

Whare are F, P, and M?

OpenStudy (anonymous):

M is the current size of resources,Pexp(at) is rate of production of new resource and Fexp(-ut) is the rate new resource deposits are being discovered

OpenStudy (anonymous):

Basically supposed to solve for the time t

OpenStudy (anonymous):

You need to solve M(t) = 0.

OpenStudy (anonymous):

So you'd have to integrate first.

OpenStudy (anonymous):

this is where i'm stuck at. M=(-1/a)Pexp(at) +P/a +F/a _(1/a)Fexp(-ut)

OpenStudy (anonymous):

Okay, what is the anti derivative of \(\exp(at)\)

OpenStudy (anonymous):

(1/a)exp(at)

OpenStudy (anonymous):

Where is the constant of integration?

OpenStudy (anonymous):

sorry. was integrated from t=0 to t=t

OpenStudy (anonymous):

Alright so what are you getting \(M(t)\)?

OpenStudy (anonymous):

M=(-1/a)Pexp(at) +P/a +F/a _(1/a)Fexp(-ut)

OpenStudy (anonymous):

How are you getting that though?

OpenStudy (anonymous):

the original equation was dM/dt=-Pexp(at)+Fexp(-ut)

OpenStudy (anonymous):

Where is +P/a coming from though?

OpenStudy (anonymous):

after the limits were put in

OpenStudy (anonymous):

from 0 to t

OpenStudy (anonymous):

I mean F/a

OpenStudy (anonymous):

Sorry. That should be F/u

OpenStudy (anonymous):

M=(-1/a)Pexp(at) +P/a +F/u -(1/u)Fexp(-ut)

OpenStudy (anonymous):

Now you would find the roots.

OpenStudy (anonymous):

Roots of which term?

OpenStudy (anonymous):

roots of t

OpenStudy (anonymous):

as in an equation of the exp(at) and exp(-ut) terms?

OpenStudy (anonymous):

Basically M = 0, and you have to find t values where that expression is 0

OpenStudy (anonymous):

The trivial case is t=0

OpenStudy (anonymous):

\[ 0=\frac Pa (1-e^{at}) + \frac Fu(1-e^{-ut}) \]

OpenStudy (anonymous):

\[ \frac Pa (1-e^{at}) = - \frac Fu(1-e^{-ut}) \]

OpenStudy (anonymous):

Hmmmm\[ -\frac{Pu}{Fa}(1-e^{at}) = 1-e^{-ut} \]

OpenStudy (anonymous):

This is the question. Obtain an analytic results that can be used for calculating the time required to consume a finite fuel resource. Currently the resource is of size, Mo, which is being produced at a rate ,P. The rate of consumption is growing exponentially at a relative rate ,Pexp(at), and new fuel resource deposits are being discovered at a rate, Fexp(-ut)

OpenStudy (anonymous):

That changes things a bit.

OpenStudy (anonymous):

So I formulated this equation dM/dt=-Pexp(at)+Fexp(-ut)

OpenStudy (anonymous):

Well, if you integrate from 0 to t you will \(\Delta M\) rather than just \(M\).

OpenStudy (anonymous):

\[ \Delta M = M-M_0 \]

OpenStudy (anonymous):

I think to find the roots in this case, we'll have to use the natural log.

OpenStudy (anonymous):

ln of -Pu/Fa=1-exp(-ut)/1-expt(at) ?

OpenStudy (anonymous):

Yeah, but we'd have to factor in the \(M_0\) as well.

OpenStudy (anonymous):

This problem is obnoxious.

OpenStudy (anonymous):

\[ M(t) = M_0 +\int_0^t\frac{dM}{dt}dt \]

OpenStudy (anonymous):

Since we want \(M(t)=0\), we have: \[ -M_0 = \int_0^t\frac{dM}{dt}dt \]

OpenStudy (anonymous):

is that a possibility?

OpenStudy (anonymous):

Yes, it's possible.